OpenStudy (anonymous):

Light travels at 186282 miles per second, and the sun is about 93 million miles from the earth. How many minutes does it take light to reach the earth from the sun? Show work

OpenStudy (anonymous):

\(\large 93,000,000 miles \cdot \frac{1second}{186,282miles} \) \(\large 93,000,000 \cancel{miles} \cdot \frac{1second}{186,282 \cancel {miles}} \) the units you're left with is seconds.... the time it takes light to travel 93 million miles

OpenStudy (anonymous):

so what is the exact answer ?